# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v1

@@ B.1 Proof of 1
where, $\eta_{n}$ is the noise vector over the $[n]$ training data. Then the posterior mean of $\widehat{\mu}$ by Gauss Markov Theorem [Johnson et al., 2002] is given by
@@ C.1 Generalization Proof
(PreDeToR risk) Suppose error stability 6.1 holds and assume loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz for all $r_{t}\in[0,B]$ and horizon $n\geq 1$. Let $\widehat{\mathrm{TF}}$ be the empirical solution of (ERM) and $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ be the covering number of the algorithm space $\mathrm{Alg}$ following Definition C.2 and C.3. Then with probability at least $1-2\delta$, the excess Multi-task learning (MTL) risk of PreDeToR-$\tau$ is bounded by
@@
We consider a meta-learning setting. Let $M$ source tasks are i.i.d. sampled from a task distribution $\mathcal{T}$, and let $\widehat{\mathrm{TF}}$ be the empirical Multitask (MTL) solution. Define $\mathcal{H}_{\text{all }}=\bigcup_{m=1}^{M}\mathcal{H}_{m}$. We drop the $\mathbf{\Theta},\mathbf{r}$ from transformer notation $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}$ as we keep the architecture fixed as in Lin et al. [2023]. Note that this transformer predicts a reward vector over the actions. To be more precise we denote the reward predicted by the transformer at round $t$ after observing history $\mathcal{H}_{m}^{t-1}$ and then sampling the action $a_{mt}$ as $\mathrm{TF}\left(\widehat{r}_{mt}(a_{mt})|\mathcal{H}_{m}^{t-1},a_{mt}\right)$. Define the training risk
@@
Using the similar steps as in Li et al. [2023] we can show that
@@
(Stability Factor) The work of Li et al. [2023] also characterizes the stability factor $K$ in 6.1 with respect to the transformer architecture. Assuming loss $\ell(\cdot,\cdot)$ is C-Lipschitz, the algorithm induced by $\mathrm{TF}(\cdot)$ obeys the stability assumption with $K=2C\left((1+\Gamma)e^{\Gamma}\right)^{L}$, where the norm of the transformer weights are upper bounded by $O(\Gamma)$ and there are $L$-layers of the transformer.
@@
(Covering Number) From Lemma 16 of Lin et al. [2023] we have the following upper bound on the covering number of the transformer class $\mathrm{TF}_{\mathbf{\Theta}}$ as
@@
where $L$ is the total number of layers of the transformer and $J$ and, $D$ denote the upper bound to the number of heads and hidden neurons in all the layers respectively. Note that this covering number holds for the specific class of transformer architecture discussed in section $2$ of [Lin et al., 2023].
@@ C.2 Generalization Error to New Task
(Dependence on $n$) In this remark, we briefly discuss why the expected excess risk for target task $\mathcal{T}$ does not depend on samples $n$. The work of Li et al. [2023] pointed out that the MTL pretraining process identifies a favorable algorithm that lies in the span of the $M$ source tasks. This is termed as inductive bias (see section 4 of Li et al. [2023]) [Soudry et al., 2018, Neyshabur et al., 2017]. Such bias would explain the lack of dependence of the expected excess transfer risk on $n$ during transfer learning. This is because given a target task $g\sim\mathcal{T}$, the $\mathrm{TF}$ can use the learnt favorable algorithm to conduct a discrete search over span of the $M$ source tasks and return the source task that best fits the new target task. Due to the discrete search space over the span of $M$ source tasks, it is not hard to see that, we need $n\propto\log(M)$ samples (which is guaranteed by the $M$ source tasks) rather than $n\propto d$ (for the linear setting).

### v2

@@ B.1 Proof of 1
where, $\eta_{n}$ is the noise vector over the $[n]$ training data. Then the posterior mean of $\widehat{\mu}$ by Gauss Markov Theorem (Johnson et al., 2002) is given by
@@ C.1 Generalization Proof
(PreDeToR risk) Suppose error stability 8.1 holds and assume loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz for all $r_{t}\in[0,B]$ and horizon $n\geq 1$. Let $\widehat{\mathrm{TF}}$ be the empirical solution of (ERM) and $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ be the covering number of the algorithm space $\mathrm{Alg}$ following Definition C.2 and C.3. Then with probability at least $1-2\delta$, the excess Multi-task learning (MTL) risk of PreDeToR-$\tau$ is bounded by
@@
We consider a meta-learning setting. Let $M$ source tasks are i.i.d. sampled from a task distribution $\mathcal{T}$, and let $\widehat{\mathrm{TF}}$ be the empirical Multitask (MTL) solution. Define $\mathcal{H}_{\text{all }}=\bigcup_{m=1}^{M}\mathcal{H}_{m}$. We drop the $\mathbf{\Theta},\mathbf{r}$ from transformer notation $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}$ as we keep the architecture fixed as in Lin et al. (2023). Note that this transformer predicts a reward vector over the actions. To be more precise we denote the reward predicted by the transformer at round $t$ after observing history $\mathcal{H}_{m}^{t-1}$ and then sampling the action $a_{mt}$ as $\mathrm{TF}\left(\widehat{r}_{mt}(a_{mt})|\mathcal{H}_{m}^{t-1},a_{mt}\right)$. Define the training risk
@@
Using the similar steps as in Li et al. (2023) we can show that
@@
(Stability Factor) The work of Li et al. (2023) also characterizes the stability factor $K$ in 8.1 with respect to the transformer architecture. Assuming loss $\ell(\cdot,\cdot)$ is C-Lipschitz, the algorithm induced by $\mathrm{TF}(\cdot)$ obeys the stability assumption with $K=2C\left((1+\Gamma)e^{\Gamma}\right)^{L}$, where the norm of the transformer weights are upper bounded by $O(\Gamma)$ and there are $L$-layers of the transformer.
@@
(Covering Number) From Lemma 16 of Lin et al. (2023) we have the following upper bound on the covering number of the transformer class $\mathrm{TF}_{\mathbf{\Theta}}$ as
@@
where $L$ is the total number of layers of the transformer and $J$ and, $D$ denote the upper bound to the number of heads and hidden neurons in all the layers respectively. Note that this covering number holds for the specific class of transformer architecture discussed in section $2$ of (Lin et al., 2023).
@@ C.2 Generalization Error to New Task
(Dependence on $n$) In this remark, we briefly discuss why the expected excess risk for target task $\mathcal{T}$ does not depend on samples $n$. The work of Li et al. (2023) pointed out that the MTL pretraining process identifies a favorable algorithm that lies in the span of the $M$ source tasks. This is termed as inductive bias (see section 4 of Li et al. (2023)) (Soudry et al., 2018; Neyshabur et al., 2017). Such bias would explain the lack of dependence of the expected excess transfer risk on $n$ during transfer learning. This is because given a target task $g\sim\mathcal{T}$, the $\mathrm{TF}$ can use the learnt favorable algorithm to conduct a discrete search over span of the $M$ source tasks and return the source task that best fits the new target task. Due to the discrete search space over the span of $M$ source tasks, it is not hard to see that, we need $n\propto\log(M)$ samples (which is guaranteed by the $M$ source tasks) rather than $n\propto d$ (for the linear setting).